TestRTP

Motoki Miura

2021 年 02 月 25 日

Contents:

第 1 章イントロダクション Intro 1 第 2 章 Arduino 3 第 3 章 Tutorial 5

3.1 The reStructuredText Cheat Sheet: Syntax Reminders . . . . . . . . . . . . . . . . . . . . . . . 5 第 4 章 Section Structure 7 第 5 章 Body Elements 9 第 6 章 Inline Markup 11 第 7 章 Directive Quick Reference 13 第 8 章 Interpreted Text Role Quick Reference 15 第 9 章 Conclusion 17 第 10 章 Indices and tables 19 参考文献 21

第 1 章

イントロダクション Intro

ここは Intro です。 images/isprime.c に、ソースコードの例を示します。

#include <stdio.h>
int isprime(int n){
int i;
for(i=2;i<n/2;i++){
  if (n%i==0) return i;
}
return 1;
}
int main(void){
int num ;  int result;
printf("input num\n") ;
scanf("%d", &num);
result = isprime(num);
if (result==1) printf("%d is prime number (素数)\n", num);
else printf("%d is not prime. %d is a divisor.\n ", num, result);
return 0;
}

リスト 1 images/isprime.c

#include <stdio.h>
int isprime(int n){
 int i;
 for(i=2;i<n/2;i++){
   if (n%i==0) return i;
 }
 return 1;
}
int main(void){
 int num ;  int result;
 printf("input num\n") ;

(次のページに続く)

\underline{TestRTP}

(前のページからの続き)

scanf("%d", &num);
result = isprime(num);
if (result==1) printf("%d is prime number (素数)\n", num);
else printf("%d is not prime. %d is a divisor.\n ", num, result);
return 0;
}

\overline{2 第 1 章イントロダクション Intro}

第 2 章

Arduino

Arduino.ino sample です。 src/sketch_feb25a.ino に、ソースコードの例を示します。強調したい場合、たとえばソースコード make は、このようになります。

リスト 1 src/sketch_feb25a.ino

    void setup() {
      // put your setup code here, to run once:
    }
    void loop() {
      // put your main code here, to run repeatedly:
    }

リスト 2 src/factorytest.ino

 1 #include "M5StickCPlus.h"
 2 #include "fft.h"
 3 #include "esp_pm.h"
 4 #include <rom/crc.h>
 5 #include <driver/i2s.h>
 6 #include <driver/rmt.h>

7

 8 #include <BLEDevice.h>
 9 #include <BLEServer.h>
10 #include <BLEUtils.h>
11 typedef struct

12 {

13  point_3d_t start_point;
14  point_3d_t end_point;
15 } line_3d_t;

16

17 typedef struct

18 {

19  double x;
20  double y;
21 } point_2d_t;

第 3 章

Tutorial

ここはチュートリアルです。

3.1 The reStructuredText Cheat Sheet: Syntax Reminders

Info See <http://docutils.sf.net/rst.html> for introductory docs. Author David Goodger <goodger@python.org> Date $Date: 2013-02-20 02:10:53 +0100 (Mi, 20. Feb 2013) $ Revision $Revision: 7612 $ Description This is a "docinfo block", or bibliographic field list

注釈: If you are reading this as HTML, please read cheatsheet.txt instead to see the input syntax examples!

第 4 章

Section Structure

Section titles are underlined or overlined & underlined.

第 5 章

Body Elements

Grid table:

Paragraphs are flush-left, separated by blank lines. Block quotes are indented.

Literal block, preceded by "::":

                                                 Indented

or:

                                                 > Quoted
 >>> print 'Doctest block'
 Doctest block

Line blocks preserve line breaks & indents. [new in 0.3.6] Useful for addresses, verse, and adornment-free lists; long lines can be wrapped with continuation lines.

Simple tables:

\underline{TestRTP}

List Type

Examples (syntax in the text source)

Bullet list • items begin with "-", "+", or "*"

Enumerated list 1. items use any variation of "1.", "A)", and "(i)" 2. also auto-enumerated

Definition list

Term is flush-left [optional classifier] Definition is indented, no blank line between

Field list field name field body

Option list

-o at least 2 spaces between option & description

Explicit Markup Examples (visible in the text source) Footnote Citation Hyperlink Target Anonymous Target

Directive ("::") Substitution Def Comment Empty Comment (".." on a line by itself, with blank lines before & after, used to separate indentation

contexts)

第 6 章

Inline Markup

emphasis; strong emphasis; interpreted text; interpreted text with role; inline literal text; standalone hyperlink, http://docutils.sourceforge.net; named reference, reStructuredText; anonymous reference; footnote reference,^{*1}; citation reference, [CIT2002]; like an inline directive; inline internal target.

*1 Manually numbered or [#] auto-numbered (even [#labelled]) or [*] auto-symbol

第 7 章

Directive Quick Reference

See <http://docutils.sf.net/docs/ref/rst/directives.html> for full info.

Directive Name Description (Docutils version added to, in [brackets])
attention  Specific admonition; also "caution", "danger", "error", "hint", "important", "note", "tip", "warning"
admonition  Generic titled admonition: ..  admonition::  By The Way
image  .. image::  picture.png; many options possible
figure  Like "image", but with optional caption and legend
topic  .. topic::  Title; like a mini section
sidebar  .. sidebar::  Title; like a mini parallel document
parsed-literal  A literal block with parsed inline markup

rubric  .. rubric::  Informal Heading

epigraph  Block quote with class="epigraph"
highlights  Block quote with class="highlights"
pull-quote  Block quote with class="pull-quote"
compound  Compound paragraphs [0.3.6]
container  Generic block-level container element [0.3.10]
table  Create a titled table [0.3.1]
list-table  Create a table from a uniform two-level bullet list [0.3.8]
csv-table  Create a table from CSV data [0.3.4]
contents  Generate a table of contents
sectnum  Automatically number sections, subsections, etc.
header, footer  Create document decorations [0.3.8]
target-notes  Create an explicit footnote for each external target
math  Mathematical notation (input in LaTeX format)
meta  HTML-specific metadata
include  Read an external reST file as if it were inline
raw  Non-reST data passed untouched to the Writer
replace  Replacement text for substitution definitions
unicode  Unicode character code conversion for substitution defs

次のページに続く \underline{TestRTP}

表 1 – 前のページからの続き

Directive Name Description (Docutils version added to, in [brackets])
date  Generates today's date; for substitution defs
class  Set a "class" attribute on the next element
role  Create a custom interpreted text role [0.3.2]
default-role  Set the default interpreted text role [0.3.10]
title  Set the metadata document title [0.3.10]

第 8 章

Interpreted Text Role Quick Reference

See <http://docutils.sf.net/docs/ref/rst/roles.html> for full info.

Role Name Description

emphasis

Equivalent to emphasis

literal

Equivalent to literal but processes backslash escapes

math

Mathematical notation (input in LaTeX format)

PEP

Reference to a numbered Python Enhancement Proposal

RFC

Reference to a numbered Internet Request For Comments

raw

For non-reST data; cannot be used directly (see docs) [0.3.6]

strong

Equivalent to strong

sub  Subscript
sup  Superscript

title Title reference (book, etc.); standard default role

17

第 9 章 Conclusion

第 10 章

Indices and tables

• genindex

• modindex

• search

参考文献

[CIT2002] A citation.